square root(3x-4) + 5 = 9
Answer by stanbon(57967)

Isolate the radical by subtracting "5" from each side.
sqrt(3x-4)=4
Square both sides to get:
3x-4 = 16
3x =20
x = 20/3
Check your answer in the original equation to make sure
it is not extraneous.
Cheers,
Stan H.